OpenStudy (anonymoussy):

Which of the following could be an example of a function with a domain [a,infinite) and a range [b,infinite) where a > 0 and b > 0?

OpenStudy (supersmart1001):

is this a multiple choice

OpenStudy (misty1212):

hi!!

OpenStudy (misty1212):

here is one that might work \[f(x)=\sqrt{x-a}+b\]

OpenStudy (misty1212):

but there are infinite choices`

OpenStudy (anonymoussy):

A. \[\sqrt{x+a}-b \] B. \[\sqrt[3]{x-b}+a\] C.\[\sqrt[3]{x+a}-b \] D.\[\sqrt{x-a}+b\]